# The global attractor of the inelastic linear Boltzmann equation in a gravity field for Maxwell molecules

### v1

@@ 1.1 The model
In the present article, we will consider a version of the inelastic Lorentz gas of Maxwell molecules evolving under the action of a uniform and fixed acceleration field, in dimension $d=2$ or $d=3$. Considering a tagged particle of such a gas, that interacts only with a set of fixed and randomly distributed scatterers, the evolution of the tagged particle is described by the linear Boltzmann equation, with a collision kernel of the form:
$B\left(\left|N\cdot\omega\right|,|v|\right)=b\left(\left|N\cdot\omega\right|\right)\hskip 8.53581pt\text{with}\hskip 8.53581pt\quad N:=\frac{v}{|v|}.$ (1.1)
(1.1) is the Maxwell molecules collision kernel. The presence of the uniform acceleration field is expressed by the presence of a drift term in the equation. We will consider the space homogeneous situation, that is, the case when the distribution function $f$ of the tagged particle is independent from the position variable, and depends only on the time variable $t$, and the velocity $v$ of the tagged particle. We will therefore consider the equation:
$\partial_{t}f(t,v)+a\cdot\partial_{v}f(t,v)=\int_{\mathbb{S}^{d-1}}\frac{1}{r}b(|\hskip 1.42262pt^{\prime}N\cdot\omega|)f(t,^{\prime}v)\mathop{}\!\mathrm{d}\omega-\left(\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)\mathop{}\!\mathrm{d}\omega\right)f(t,v),$ (1.2)
where we used the notation $\hskip 1.42262pt{}^{\prime}N=^{\prime}v/|^{\prime}v|$, $a\in\mathbb{R}^{d}$ is the acceleration field, and ${}^{\prime}v$ is the pre-collisional velocity of $v$, that is, such that a particle colliding with a scatterer with velocity ${}^{\prime}v$ and angular parameter $\omega$ is reflected with the velocity $v$. Finally, we assume that the inelastic collisions take place according to the following collision law:
$v=^{\prime}v-(1+r)(^{\prime}v\cdot\omega)\omega\hskip 8.53581pt\text{or equivalently}\hskip 8.53581pt^{\prime}v=v-(1+1/r)(v\cdot\omega)\omega,$ (1.3)
@@
$=T^{\mu}(t)\int_{\mathbb{S}^{d-1}}\frac{1}{r}b(|\hskip 0.7113pt^{\prime}N\cdot\omega|)f(t,^{\prime}v)\mathop{}\!\mathrm{d}\omega-T^{\mu}(t)\left(\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)\mathop{}\!\mathrm{d}\omega\right)f(t,v)$ (1.4)
@@
[…]
@@ 1.2 Lorentz gas and granular material: a brief review of the literature
Introduced by H. A. Lorentz in 1905, the Lorentz gas model has been originally proposed to analyze electron transport in metals. It describes the dynamics of non-interacting particles moving through a fixed random configuration of heavy scatterers. The interaction between the Lorentz particle and the scatterers is characterized by a finite-range central potential. Despite its simplicity, this model is profoundly important, since it is one of the few sources of exact results in kinetic theory, showcasing how microscopic reversibility can be reconciled with macroscopic irreversibility. Specifically, it allows for the demonstration of the validity of linear kinetic equations and, subsequently, diffusion equations under appropriate scaling limits. The first rigorous proof of the linear Boltzmann equation’s validity was established in [37] within the Boltzmann–Grad limit. For further related findings and developments, we refer to references [4, 19, 49, 61, 62, 72, 73]. The rigorous derivation of hydrodynamic equations, specifically the heat equation, from the random Lorentz gas remains a challenging and unresolved problem. However, it is noteworthy that the rigorous validity of the diffusive limit has been established by Bunimovich and Sinai (see [24]) for periodically distributed scatterers. In the random case, the diffusion equation can be derived by relying on the kinetic approximation of the microscopic dynamics. Specifically, the heat equation can be obtained from the random Lorentz gas employing the kinetic equation that emerges in an appropriate kinetic limit as a bridge. This approach has been used in [18], [3], [4], and also in [34]. Additionally, we refer to [49] for an alternative approach in this direction.
In most of the mathematical studies of the linear Boltzmann equation, it has been assumed that additional transport terms describing the effect of possible external fields are absent. Notably, the presence of an external field significantly influences the derivation of the linear Boltzmann equation in low-density regimes, as well as the properties of its solutions. More precisely, in [13] and later in [14] (see also [44] where the model has been studied numerically), it has been shown that the motion of a test particle in the Euclidean space $\mathbb{R}^{2}$ with a Poisson distribution of hard disk scatterers and a uniform, constant magnetic field perpendicular to the plane formally leads to the generalized Boltzmann equation with memory terms. A rigorous derivation of this linear Boltzmann equation has been recently obtained in [63]. We also refer to [50], where linear kinetic equations with a magnetic transport term have been derived, but the non-Markovian behavior of the limit process disappears. On the other hand we also mention [20] where the well-posedness of a slightly different linear Boltzmann equation with external field, specifically a space periodic electric field, has been studied in the time-stationary regime.
@@
The inelastic collision model.
Turning to the case of dissipative gases, we first mention that considering inelastic collisions between the elementary components of a fluid is motivated by many physical situation: snow, sand or interstellar dust may be modelled by a very large number of particles that dissipate kinetic energy along collisions. Such a model has countless applications, in theoretical physics as well as in industry (see in particular [42]).
The question of modelling inelastic collisions is delicate. It may be performed by studying the deformation of colliding objects and the contact force that is generated by such a deformation, as it was first done by Hertz in the elastic case [40], and later adaptated by Brilliantov et al. [22] in the inelastic case. Determining the contact force enables then to deduce the post-collisional relative velocity (right when the particles cease to be in contact) from the pre-collisional relative velocity (at the time the particles enter in contact). Since the time during which the particles are in contact and deformed is typically much smaller than the time scale used to describe the motion of the whole system, it is therefore usual in kinetic theory to consider that collisions take place instantaneously.
One major simplification of the previous description consists in considering interactions without friction, so that only the normal component $v\cdot\omega$ (that is, along the line of contact between the particles, orientated by $\omega\in\mathbb{S}^{d-1}$) of the relative velocity $v$ is modified during the collisions, so that the tangential component remains unchanged. As a consequence, to completely determine the law of collision in the frictionless case, it is enough to prescribe a single real constant, the so-called restitution coefficient (denoted by $r$), defined as the absolute value of the ratio between the normal component $v^{\prime}\cdot\omega$ of the post-collisional relative velocity $v^{\prime}$ and the normal component $v\cdot\omega$ of the pre-collisional relative velocity $v$:
$r=\left|\frac{v^{\prime}\cdot\omega}{v\cdot\omega}\right|\cdotp$ (1.6)
In the physics literature, based on experiments (see [23] and the references therein) as well as on theoretical considerations ([69], [67], [23]), it is well-known that $r$ depends on the norm of the relative velocity of the colliding particles, and tends to $1$ when the relative velocity becomes small ($r=1$ corresponds to an elastic collision). Particle systems with variable restitution coefficients are called viscoelastic particle systems. Nevertheless, assuming that the restitution coefficient $r\in[0,1]$ is a fixed constant yields a much simpler model, which is still a good approximation of real particles (see the discussion in [23]).
We finally mention that more accurate descriptions taking into account friction exist, relying on another restitution coefficient used to model the change of the tangential component of the relative velocity during the collisions. The reader may refer to [6], [70], and again to [23] and the references therein. In the present article, we will focus on the frictionless model, and although we will briefly discuss the literature including also the case of viscoelastic particles (variable restitution coefficient) in the introductory Section 1.3, we will consider only with fixed (normal) restitution coefficient in our investigations.
@@
Fluids or solids composed with a large number of particles that interact inelastically are called granular media. Such systems present a wide range of very interesting behaviours, and exhibit differences of fundamental nature with respect to conservative fluids, the evolution of which is governed by the classical, elastic Boltzmann equation. For a brief introduction to the peculiarities of the granular media, the reader may consult the surveys [42] and [43]. Let us present two of such specificities, which are maybe the most prominent.
Within granular media, the dissipation of kinetic energy along collisions induces a decay of the temperature in the case when there is no external source of energy injected in the system. In many cases, it is possible to prove that the temperature decays according to a power law: this is the celebrated Haff’s law [39]. Observe that the algebraic exponent of such a decay depends on the properties of the restitution coefficient (see [23]).
@@
In addition, in the case of a fixed restitution coefficient that is small enough, inelastic particle systems exhibit inelastic collapse: infinitely many collisions take place in finite time. At the level of the kinetic equation, this phenomenon can be identified by the convergence in finite time of the solutions towards Dirac masses, which represents a brutal concentration of the velocity profile of the distribution function. A vast literature exists about this phenomenon: see [54], [23], [66], [30] and the references therein, and [31] for the $d$-dimensional case with $d\geq 2$. Concerning the one-dimensional case (which is not only a toy model but turns out to be of fundamental importance according to the observations of [54], as well as a challenging mathematical problem), the reader may refer to [10], and the more recent [32] and the references therein.
On the one hand, the inelastic collapse phenomenon can be seen as an extreme case of clusterization. Such a phenomenon constitutes a major obstruction to perform numerical simulations, and to provide a rigorous derivation of inelastic kinetic equations (no analog of Lanford’s theorem [45] exists for such models). On the other hand, it is still unclear if such a phenomenon exists for models different from the particular, not completely physical case of fixed restitution coefficient, nor if the dynamics of the particle system is globally well-posed, even when a collapse occurs.
@@ 1.3 Mathematical theory of the inelastic Boltzmann equation
In this section, we briefly review the state of the art concerning the inelastic Boltzmann equation. In order to insert the results of the present article in the broad context of the inelastic Boltzmann equation(s), we will consider the non-linear case in a first time, corresponding to a gas of particles that interact with each other. In a second time, we will turn to the linear case, corresponding to particle systems that do not interact with each other, but only with a background, that is either composed with another gas at thermal equilibrium (Rayleigh gas) or with fixed obstacle (Lorentz gas).
In order to establish inelastic versions of the Boltzmann equation, the starting point is the weak formulation (for instance, see [25]). The weak formulation is a consequence of the formal derivation of the Boltzmann equation, which is performed via a careful counting of the different ways to modify the number of particles of the system with a velocity that belongs to the infinitesimal volume $v+\mathop{}\!\mathrm{d}v$ of the phase space. In the linear case, which is the main focus of the present article, the equation reads (for a general collision kernel):
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\int_{\mathbb{R}^{d}}\varphi(v)f(t,v)\mathop{}\!\mathrm{d}v=\int_{\mathbb{R}^{d}}\int_{\mathbb{S}^{d-1}}B\left(|N\cdot\omega|,|v|\right)f(t,v)\left[\varphi(v^{\prime})-\varphi(v)\right]\mathop{}\!\mathrm{d}\omega\mathop{}\!\mathrm{d}v,$ (1.7)
for any test function $\varphi$, and where $v^{\prime}$ is the post-collisional velocity given by the expression:
$v^{\prime}=v-\left(1+r\right)(v\cdot\omega)\omega.$ (1.8)
We denote by $c$ the scattering mapping (1.8) that to $v$ associates $v^{\prime}$, and by $J_{c}$ its Jacobian determinant:
$J_{c}=J(v^{\prime})=J(c(v)).$ (1.9)
In the case of a fixed restitution coefficient $r$ we have:
$\left|J_{c}(v)\right|=r.$ (1.10)
The strong form of the inelastic, spatially homogeneous, linear Boltzmann equation with a general collision kernel $B$, is the following:
$\partial_{t}f(t,v)$ $=\int_{\mathbb{S}^{d-1}}B\left(|\hskip 1.42262pt^{\prime}N\cdot\omega|,|v|\right)\frac{1}{|Jc(^{\prime}v)|}f(^{\prime}v)\mathop{}\!\mathrm{d}\omega-f(v)\int_{\mathbb{S}^{d-1}}B(|N\cdot\omega|,|v|)\mathop{}\!\mathrm{d}\omega,$ (1.11)
where ${}^{\prime}v$ is the pre-collisional velocity obtained with inelastic collisions:
${}^{\prime}v=v-\left(1+1/r\right)(v\cdot\omega)\omega.$ (1.12)
Contrary to the case of the elastic Boltzmann equation, a non-trivial Jacobian $J_{c}$ appears in the gain term, when the collision operator is written in the strong form. This Jacobian corresponds to the change of variables from post- to pre-collisional velocities. In the elastic case we have $|J_{c}|=1$.
@@
In the non-linear case, the strong form of the inelastic Boltzmann equation reads in general:
$\partial_{t}f(t,x,v)+v\cdot\partial_{x}f(t,x,v)=\int_{\mathbb{S}^{d-1}}\int_{\mathbb{R}^{d}}\Bigg{[}$ $\frac{B(|\hskip 0.7113pt^{\prime}v-\hskip 0.7113pt^{\prime}v_{*}|,\left|\hskip 1.42262pt{}^{\prime}N\cdot\omega\right|)}{|Jc(^{\prime}v,^{\prime}v_{*})|}f(t,x,^{\prime}v)f(t,x,^{\prime}v_{*})$
$\hskip 28.45274pt-B(|v-v_{*}|,\left|N\cdot\omega\right|)f(t,x,v)f(t,x,v_{*})\Bigg{]}\mathop{}\!\mathrm{d}v_{*}\mathop{}\!\mathrm{d}\omega,$ (1.13)
where $\hskip 0.7113pt{}^{\prime}v$, $\hskip 0.7113pt{}^{\prime}v_{*}$ denote the pre-collisional velocities, defined as:
$\hskip 0.7113pt^{\prime}v=v-\frac{(1+1/r)}{2}(v-v_{*})\cdot\omega\omega,\hskip 14.22636pt\hskip 0.7113pt^{\prime}v_{*}=v_{*}+\frac{(1+1/r)}{2}(v-v_{*})\cdot\omega\omega,$ (1.14)
or equivalently, $(v,v_{*})=c(\hskip 0.7113pt^{\prime}v,\hskip 0.7113pt^{\prime}v_{*})$, with $c(v,v_{*})=\left(c_{1}(v,v_{*}),c_{2}(v,v_{*})\right)$, and
$c_{1}(v,v_{*})=v-\frac{(1+r)}{2}(v-v_{*})\cdot\omega\omega,\hskip 14.22636ptc_{2}(v,v_{*})=v_{*}+\frac{(1+r)}{2}(v-v_{*})\cdot\omega\omega,$ (1.15)
with $\omega\in\mathbb{S}^{d-1}$ the angular parameter (the orientation of the line of contact of two colliding particles in the case of hard spheres), $N$, $\hskip 1.42262pt{}^{\prime}N$ denoting:
$N=\frac{v-v_{*}}{|v-v_{*}|},\hskip 14.22636pt\hskip 1.42262pt^{\prime}N=\frac{\hskip 0.7113pt{}^{\prime}v-\hskip 0.7113pt^{\prime}v_{*}}{|\hskip 0.7113pt^{\prime}v-\hskip 0.7113pt^{\prime}v_{*}|},$ (1.16)
and $Jc$ being the Jacobian of the mapping $(v,v_{*})\mapsto c(v,v_{*})$ (considered at $\omega$ fixed).
Going back to the weak form, we recover directly the conservation of the total mass $\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}f(t,x,v)\mathop{}\!\mathrm{d}v\mathop{}\!\mathrm{d}x$ and the momentum $\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}vf(t,x,v)\mathop{}\!\mathrm{d}v\mathop{}\!\mathrm{d}x$.
@@
We refer to [76] for a very complete survey about the mathematical state of the art concerning (1.3), and to the more recent [27], which addresses also the question of the numerical simulations related to (1.3).
Contrary to the elastic Boltzmann equation, the dissipative character of the collisions causes the solutions of (1.3) to develop singularities (concentrations of the velocity profile, explosion of the gradient $\partial_{x}f$…), and the trend to create clusters prevents to consider linearization around an homogeneous solution. For these reasons, the mathematical theory of (1.3) is still quite incomplete, and many results are obtained only in particular cases, or for related, simpler models.
Rigorous results were first obtained in the one-dimensional case (observe that, contrary to the elastic Boltzmann equation, the one-dimensional case yields a non-trivial collision operator), relying on the so-called quasi-elastic limit: a Boltzmann-like equation is derived using the BBGKY hierarchy for a system of $N$ inelastic hard spheres ([7], see also [53] and [74] for other derivations), assuming that the restitution coefficient $r$ is a function $r(N)$ of the number of particles $N$, such that $r(N)\rightarrow 1$, as $N\rightarrow+\infty$, which prevents in particular the inelastic collapse to take place ([10]). The kinetic equation that is obtained is called the non-linear friction equation ([74]), and can be written as:
$\partial_{t}f(t,x,v)+v\partial_{x}f(t,x,v)=-\partial_{v}\left[\int_{\mathbb{R}}(v_{*}-v)|v_{*}-v|f(t,x,v_{*})\mathop{}\!\mathrm{d}v_{*}\right].$ (1.17)
In the space-homogeneous case, the solutions can be studied completely ([7]): the equation is globally well-posed for any initial datum that is a Borel probability measure with finite moment of order $2$ and finite entropy, and if the initial datum admits a $\mathcal{C}^{k}$ density (with respect to the Lebesgue measure), then so does the solution. In addition, the asymptotic behaviour of the solutions can be precisely described: all the solutions converge towards a Dirac mass, and up to consider solutions obtained from $L^{1}$ initial data, after a proper rescaling it is possible to show that the rescaled solutions converge towards the sum of Dirac masses $\frac{1}{2}\delta_{-1/2}+\frac{1}{2}\delta_{1/2}$. Note that this last result reminds the spontaneous onset of shear flows in granular media observed in [55] when the restitution coefficient is chosen in between the elastic kinetic regime, and the typically inelastic regime of clusterization.
Later, Toscani [74] studied the one-dimensional inelastic Boltzmann equation with the hard sphere collision kernel, with general restitution coefficients:
$\partial_{t}f+v\partial_{x}f=\int_{\mathbb{R}}\int_{0}^{+\infty}\beta(\theta)\left[\frac{|\hskip 1.42262pt^{\prime}v-\hskip 0.28453pt^{\prime}v_{*}|}{J_{c}}f(t,x,^{\prime}v)f(t,x,^{\prime}v_{*})-|v-v_{*}|f(v)f(v_{*})\right]\mathop{}\!\mathrm{d}\theta\mathop{}\!\mathrm{d}v_{*},$ (1.18)
where the restitution coefficient $r$, involved in the scattering formulas (1.14) and (1.15), is given by:
$r=r(\theta,|v-v_{*}|)=\frac{1}{1+\theta|v-v_{*}|^{\gamma}},$ (1.19)
with $\gamma\in\mathbb{R}$ a fixed constant and $\theta\geq 0$. This model covers therefore a particular case of viscoelastic particles. $\theta$ measures the degree of inelasticity ($\theta=0$ corresponds to an elastic collision, while $\theta=+\infty$ corresponds to a perfectly inelastic collision), and $\beta(\theta)$ is the probability that a collision with degree of inelasticity $\theta$ takes place. In particular, considering $\gamma=0$ and $\beta=\delta_{(1-r_{0})/r_{0}}$ with $r_{0}\in]0,1]$, one recovers the model with fixed restitution coefficient $r$, equal to $r_{0}$, so that (1.18) becomes:
$\partial_{t}f+v\partial_{x}f=\int_{\mathbb{R}}|v-v_{*}|\left[\frac{1}{r^{2}}f(t,x,^{\prime}v)f(t,x,^{\prime}v_{*})-f(v)f(v_{*})\right]\mathop{}\!\mathrm{d}v_{*}.$ (1.20)
In [74], it is in particular proved that the space-homogeneous version of (1.18) is well-posed, in the same spaces as for (1.17). Then, Benedetto, Caglioti and Pulvirenti obtained results in the non-homogeneous case. In [8], it is proved that (1.20) (corresponding to the case of a fixed restitution coefficient) is well-posed close to the vacuum. More specifically here, it is assumed that the initial datum $f_{0}$ belongs to $L^{1}(\mathbb{R}^{2})\cap L^{\infty}(\mathbb{R}^{2})$, vanishes for large enough velocities, and such that $\|f_{0}\|_{L^{1}}$ is small enough. Finally in [11], considering the case $\gamma>0$ (and $\beta$ being a Dirac mass), the well-posedness of (1.18) is proved without assuming anything on the $L^{1}$ norm of $f_{0}$ (but $f_{0}$ is still assumed to be vanishing for $|v|$ large).
@@
Later, many authors turned to the case of higher dimension, motivated to investigate a conjecture due to Ernst and Brito [35]. Such a conjecture postulated that, while the temperature of a granular gas decays according to the Haff’s law, the velocity profile $f(t,v)$ should converge towards a self-similar solution of the inelastic Boltzmann equation:
@@
where $\alpha,\beta:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}^{*}$ are two scaling functions. In addition, the conjecture postulated that the self-similar profile $m$, contrary to the Maxwellian equilibria of the elastic Boltzmann equation, should present overpopulated high energy tails, in the sense that there exists some $20<b<$ 2 and $a>0$ such that:
$m(|v|)\sim e^{-a|v|^{b}}\hskip 8.53581pt\text{as}\ |v|\rightarrow+\infty.$ (1.22)
[…]
In order to restore also the existence of steady states, variations of the space-homogeneous version of (1.3) for Maxwell molecules were considered. For instance, it is possible to consider the inelastic particles in a thermal bath, reinjecting the energy that is lost during the collisions (first in [9], then, for instance, in [26]). Another effect competing with the inelastic collisions that was considered is the presence of a shear [28].
Finally, important progresses were completed in the series of articles [58], [59] and [60], in which the space-homogeneous, inelastic Boltzmann equation (1.3) is considered for the hard sphere collision kernel, which corresponds better to a physical model than the Maxwell molecule kernel. In particular, the Cauchy theory is established in a quite general setting, and the Haff’s law as well as the Ernst-Brito conjecture are proved for restitution coefficients close enough to $1$. We mention also [2], in which the Haff’s law is proved for viscoelastic particles, with optimal lower and upper bounds.
To conclude this discussion concerning the non-linear inelastic Boltzmann equation, we emphasize that, to the best of our knowledge, the only reference that deals with the spatially non-homogeneous version of (1.3) (in general dimension) is [1], where the well-posedness of the solutions is obtained only close to the vacuum.

### v2

@@ 1.1 The model
In the present article, we will consider an inelastic Lorentz gas with Maxwell molecule interactions under the action of a uniform and fixed acceleration field, in dimension $d=2$ or $d=3$. The evolution of the tagged particle interacting inelastically with a set of fixed and randomly distributed scatterers is described by the inelastic linear Boltzmann equation, with collision kernel of the form:
$B\left(\left|N\cdot\omega\right|,|v|\right)=b\left(\left|N\cdot\omega\right|\right)\hskip 8.53581pt\text{with}\hskip 8.53581pt\quad N:=\frac{v}{|v|},$ (1.1)
which is the Maxwell molecules collision kernel. The presence of the uniform acceleration field is expressed by the presence of a drift term in the equation. We will consider the space homogeneous situation, that is, the case when the distribution function $f$ of the tagged particle is independent from the position variable, and depends only on the time variable $t$, and the velocity $v$ of the tagged particle. We will therefore consider the equation:
$\partial_{t}f(t,v)+a\cdot\partial_{v}f(t,v)=\int_{\mathbb{S}^{d-1}}\frac{1}{r}b(|\hskip 1.42262pt^{\prime}\hskip-1.42262ptN\cdot\omega|)f(t,^{\prime}\hskip-2.84526ptv)\mathop{}\!\mathrm{d}\omega-\left(\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)\mathop{}\!\mathrm{d}\omega\right)f(t,v),$ (1.2)
where we used the notation $\hskip 1.42262pt{}^{\prime}\hskip-1.42262ptN=\hskip 0.0pt^{\prime}\hskip-1.42262ptv/|^{\prime}\hskip-1.42262ptv|$, $a\in\mathbb{R}^{d}$ is the acceleration field, and $^{\prime}\hskip-1.42262ptv$ is the pre-collisional velocity of $v$, that is, such that a particle colliding with a scatterer with velocity $^{\prime}\hskip-1.42262ptv$ and angular parameter $\omega$ is reflected with the velocity $v$. Finally, we assume that the inelastic collisions take place according to the following collision law:
$v=\hskip 0.0pt^{\prime}\hskip-1.42262ptv-(1+r)(^{\prime}\hskip-1.42262ptv\cdot\omega)\omega\hskip 8.53581pt\text{or equivalently}\hskip 8.53581pt^{\prime}\hskip-1.42262ptv=v-(1+1/r)(v\cdot\omega)\omega,$ (1.3)
@@
$=T^{\mu}(t)\int_{\mathbb{S}^{d-1}}\frac{1}{r}b(|\hskip 1.42262pt^{\prime}\hskip-2.13394ptN\cdot\omega|)f(t,^{\prime}\hskip-2.84526ptv)\mathop{}\!\mathrm{d}\omega-T^{\mu}(t)\left(\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)\mathop{}\!\mathrm{d}\omega\right)f(t,v)$ (1.4)
@@
$T(t)=\frac{1}{2}\int_{\mathbb{R}^{d}}|v|^{2}f(t,v)\mathop{}\!\mathrm{d}v\hskip 5.69054pt\text{and}\hskip 5.69054pt\mu\geq 0.$ (1.5)
@@ 1.2 Lorentz gas and granular material: a brief review of the literature
The Lorentz gas model, introduced by H. A. Lorentz in 1905 to study electron transport in metals, describes the dynamics of non-interacting particles moving through a fixed random configuration of heavy scatterers. The interaction between the Lorentz particle and the scatterers is characterized by a finite-range central potential. Despite its simplicity, it represents a yet profoundly significant model providing key insights on how microscopic reversibility can be reconciled with macroscopic irreversibility. In fact, from this model, one can obtain, under suitable scaling limits, linear kinetic equations ([31, 17, 42, 52, 53, 59, 60]), and, from these, one can derive diffusion equations in longer time scales (cf. [16, 5, 6, 42]).
In most of the mathematical studies of the linear Boltzmann equation, it has been assumed that additional transport terms describing the effect of possible external fields are absent. However, the presence of external fields significantly affects its derivation in low-density regimes, as well as the properties of its solutions. More precisely, in [12] and later in [13] (see also [37] where the model has been studied numerically), it has been shown that the motion of a test particle in $\mathbb{R}^{2}$ with a Poisson distribution of hard disk scatterers and a uniform, constant magnetic field perpendicular to the plane formally leads to the generalized Boltzmann equation with memory terms. A rigorous derivation of this equation has been recently obtained in [54]. We also refer to [43], where linear kinetic equations with a magnetic transport term have been derived. On the other hand we also mention [18] where the well-posedness of a slightly different linear Boltzmann equation with external field, specifically a space periodic electric field, has been studied in the time-stationary regime.
@@
Turning to the case of dissipative gases, we first mention that considering inelastic collisions between the elementary components of a fluid is motivated by many physical situation: snow, sand or interstellar dust may be modelled by a very large number of particles that dissipate kinetic energy during collisions. Such a model has countless applications, in theoretical physics as well as in industry (see in particular [35]). Systems composed of a large number of particles that interact inelastically, the so-called granular media, present a wide range of very interesting behaviours, and exhibit differences of fundamental nature with respect to conservative fluids, the evolution of which is governed in the low density limit by the classical, elastic Boltzmann equation. For a brief introduction to the peculiarities of the granular media, the reader may consult the surveys [35] and [36]. Let us present two of such specificities, which are maybe the most prominent.
Within granular media, the dissipation of kinetic energy during collisions induces a decay of the temperature in the case when there is no external source that is injecting energy in the system. In many cases, it is possible to prove that the temperature decays according to a power law: this is the celebrated Haff’s law [33]. Observe that the algebraic exponent of such a decay depends on the properties of the restitution coefficient (see [20]).
@@
In addition, in the case of a fixed restitution coefficient that is small enough, inelastic particle systems exhibit inelastic collapse: infinitely many collisions take place in finite time. At the level of the kinetic equation, this phenomenon can be identified by the convergence in finite time of the solutions towards Dirac masses, which represents a brutal concentration of the velocity profile of the distribution function. A vast literature exists about this phenomenon: see [46], [10], [20], [56], [25], [27], [26], [28] and the references therein.
On the one hand, the inelastic collapse phenomenon can be seen as an extreme case of clusterization. Such a phenomenon constitutes a major obstruction to perform numerical simulations, and to provide a rigorous derivation of inelastic kinetic equations (no analog of Lanford’s theorem [38] exists for such models). On the other hand, it is still unclear if such a phenomenon exists for models different from the particular case of fixed restitution coefficient, nor if the dynamics of the particle system is globally well-posed, even when a collapse occurs (see [25], [26]).
@@ 1.3 Mathematical theory of the inelastic Boltzmann equation
In this section, we briefly review the state of the art concerning the inelastic Boltzmann equation. Contrary to the elastic Boltzmann equation, the dissipative character of the collisions causes the solutions to develop singularities (concentrations of the velocity profile, explosion of the gradient $\partial_{x}f$…), and the trend to create clusters prevents to consider linearization around an homogeneous solution. For these reasons, the mathematical theory of the inelastic Boltzmann equation is still quite incomplete, and many results are obtained only in particular cases, or for related, simpler models.
@@
We refer to [62] for a very complete survey about the mathematical state of the art concerning the non-linear version of the equation, and to the more recent [23], which addresses also the question of the numerical simulations.
@@
Rigorous results were first obtained in the one-dimensional case ([7], [8], [61]), before that the case of higher dimension was studied by many authors. This research was motivated by from a conjecture due to Ernst and Brito [29], postulating that, while the temperature of a granular gas decays according to the Haff’s law, the velocity profile $f(t,v)$ should converge towards a self-similar solution of the inelastic Boltzmann equation:
@@
where $\alpha,\beta:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}^{*}$ are two scaling functions, and such that the self-similar profile $m$ should present overpopulated high energy tails: $m(|v|)\sim e^{-a|v|^{b}}$ as $|v|\rightarrow+\infty$ for some $0<b<2$ and $a>0$.
At the same time, a simplified version of the inelastic Boltzmann equation was introduced, corresponding to Maxwell molecule interaction (the assumption (1.1) on the collision kernel). In such a case, the evolution of the moments can be determined. Nevertheless, as observed for instance in [62], such a model would lead to an exponential decay of the temperature, which does not correspond to the algebraic decay prescribed by the Haff’s law. For this reason, several authors ([22], [14]) considered a rehomogeneization factor $T^{\mu}(t)$ in front of the collision operator (as in (1.1)) in order to restore the homogeneity of the collision operator, and to recover an ad hoc Haff’s law.
In order to restore also the existence of steady states to the inelastic Boltzmann equation, variations of the space-homogeneous version for Maxwell molecules were considered. For instance, it is possible to consider the inelastic particles in a thermal bath, reinjecting the energy that is lost during the collisions (first in [9], then, for instance, in [22]). Another effect competing with the inelastic collisions that was considered is the presence of a shear [24].
Finally, important progresses were completed in the series of articles [49], [50] and [51], in which the space-homogeneous, inelastic Boltzmann equation is considered for the hard sphere collision kernel, which corresponds better to a physical model than the Maxwell molecule kernel. In particular, the Cauchy theory is established in a quite general setting, and the Haff’s law as well as the Ernst-Brito conjecture are proved for restitution coefficients close enough to $1$. We mention also [3], in which the Haff’s law is proved for viscoelastic particles, with optimal lower and upper bounds.
To conclude this discussion concerning the non-linear inelastic Boltzmann equation, we emphasize that, to the best of our knowledge, the only references that deal with the spatially non-homogeneous version of (in general dimension) are [2] and [4], where the well-posedness of the solutions is obtained only close to the vacuum in [2], and close to thermal equilibrium, for restitution coefficients $r$ close enough to $1$ (that is, close enough to the elastic case) in [4].